OpenStudy (anonymous):

please help.....What is the value of x so that the line segment with endpoints W(x, −2) and X(5, −4) is parallel to the line segment with endpoints Y(2, 2) and Z(5, 6)? x = 6 1/2 x = 6 x = 3 1/2 x = 7

OpenStudy (sweetburger):

parallel lines have equal slopes

OpenStudy (sweetburger):

Given you can find the slope of one of the lines you can set the slope you find equal to the equation to find the slope of the other line to produce an x value.

OpenStudy (jdoe0001):

http://openstudy.com/updates/56d8d403e4b07475bd9ae984 and so.. have you covered slopes then? if you haven't, this may be a good time to do so firstly

OpenStudy (anonymous):

Let's ignore the multiple choices, and focus on the goal. Find the value of x so the line segment with endpoints (x, −2) and (5, −4) is parallel to the line segment with endpoints (2, 2) and (5, 6) The second line has a slope of 4/3. To be parallel to each other, the first line's slope is -2 / 5-x or 4/ 2x-10 so 2x-10=3 and x = 13/2 = 6 1/2

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Finding the slope of the second line with endpoints (2, 2) and (5, 6) : slope, or 'm', is a ratio of the y change and the x change between two points. so m = Δy / Δx *Δ is a symbol which represents "change in" or "change between". in our case, we're finding Δy which is the "change in y", or Δx "the change in x" Find the slope m between ponts (2, 2) and (5, 6). m = Δy / Δx Δy = from 2 to 6 = 4 Δx = from 2 to 5 = 3 Form the ratio m = Δy / Δx and we find m = 4 / 3

OpenStudy (anonymous):

Finding the slope of the first line with endpoints (x, -2) and (5, -4). slope m is the ratio of Δy and Δx. Δy = from -2 to -4 = -2 Δx = from x to 5 = 5 - x our slope m = Δy /Δx = -2 /[5 - x] *lines are parallel when their slopes equal each other. This means our first slope 4 /3 equals our second slope. 4 /3 = -2 /[5 - x] *This would be nicer if our numerators equaled each other. Multiplying the right part of the equation by -2 / -2 means : 4 /3 = 4 /-2[5 - x] Equating the denominators of each side of the equation means : 3 = -2[5 - x] *Remember to distribute the -2 coefficient lying in front of the [5 - x] 3 = -10 + 2x Solving this equation means : 13 = 2x x = 13 /2 = 6 and 1 /2

OpenStudy (anonymous):

okay got it, that helps thank you!